1

Problem sheet 9, 2005, Dec. 5th  MT441 Channels

Good further problems for exam preparation purposes can be found on the course’s webpage and at Prof. Cover’s webpage. http://www.stanford.edu/class/ee376a/ Ignore problems on material that is not covered in this course...

Ex. 1 Read the description of the Lempel Ziv algorithm in McKay’s book. Describe, in your own words the method, and briefly discuss its performance, advantages/disadvantages. How good would it be for large examples? Can the basic method be improved?

Ex. 2 Encode the string using the Lempel Ziv method.

000000000000100000000000.

Ex. 3 Decode the string

00101011101100100100011010101000011,

which was encoded by the basic Lempel Ziv encoding in McKay’s notation.

If you return it by Monday, I will mark it by Wednesday (last lecture of this course).

If you want to see any particular section revised in the last lecture, let me know (the sooner the better). My web page contains a collection of related material. http://www.ma.rhul.ac.uk/∼elsholtz/WWW/lectures/0506mt441/lecture.html